# Soundscape Captioning using Sound Affective Quality Network and Large Language Model

### v2

@@ IV-C Acoustic model results and analysis
6) RQ6: Does SoundAQnet differentiate between acoustic scenes (ASs) in affective quality (AQ) predictions?
Figure 7: The scatter plot, density distribution curves, and marginal distributions for the 3 scenes involved are from SoundAQnet’s predictions on the test set.
@@
V Human expert evaluation
@@
V-A Experimental design for quality assessment
The study employs a within-subjects design to evaluate soundscape captions from SoundSCaper and human experts. The sample size calculation is performed using G*Power [51]. The results of the calculation indicated that a sample size of 30 audio samples with $\alpha=0.05$ and an assumed Effect Size of 0.5 for the Wilcoxon signed rank test achieved the pre-statistical power of 83.3%. Thus, the evaluation contains 60 audio clips from two distinct datasets. Dataset 1 (D1) contains 30 randomly selected samples with the same sound pressure levels (SPLs) from this paper’s test set. Dataset 2 (D2) has 30 samples randomly selected from 5 external, i.e., model-never-seen, audio scene datasets, which are DCASE 2018/2019 [52], ISD [53], LITIS-Rouen [54] and road environment datasets [55]. This paper’s training set contains 3 types of acoustic scenes, so recordings related to the 3 AS labels in this paper are selected from the 5 external datasets. Finally, the total duration of the D2 candidate data pool is about 1177 hours. Audio clips in D2 vary from 10 to 30 seconds with various SPLs without any limitations. Hence, D2 is mainly used to test the generalization performance of the SoundSCaper system.
V-A1 Human expert annotations
@@
V-A2 SoundSCaper captions
As described in Section III-B, SoundSCaper automatically generates target descriptions.
Finally, 120 soundscape captions are evaluated, 60 of which are derived from the proposed SoundSCaper system, and the remaining 60 are annotated by the two experienced soundscape experts. These captions are randomized. A jury of 16 audio/soundscape experts evaluated these captions. Human assessment instructions, assessment data, and expert assessment statistics’ metadata are all public on the homepage.
V-A3 Ethic permission
As the primary institution of this paper, Ghent University (UGent) adheres to a strict code of ethics and complies with the General Data Protection Regulation (GDPR). Based on a self-assessment of the study’s risks, ethical approval for the research in the paper was obtained from the Faculty of Engineering and Architecture of UGent.
V-B Soundscape caption evaluation metrics
Inspired by [56], we introduce the Transparent Human Benchmark for Soundscapes (THumBS) as a metric for soundscape captions. This indicator consists of precision, recall, and other three types of penalty items targeting specific defects.
V-B1 Precision and recall $\in[1,5]$
@@ V-B3 THumBS score
Due to space limitations, we fully explain these metrics in the participant instructions presented on the homepage.
V-C Soundscape caption evaluation result and analysis
1) RQ1: How does it differ between the soundscape captions generated by SoundSCaper and those by human experts?
TABLE VI: Comparison of soundscape caption quality from human expert (H) and SoundSCaper (S) on datasets D1 and D2.
@@ V-C Soundscape caption evaluation result and analysis
1 H 3.84$\pm$0.30 3.93$\pm$0.21 -0.10$\pm$0.07 -0.14$\pm$0.12 -0.22$\pm$0.12 3.43$\pm$0.35
S 3.79$\pm$0.39 3.86$\pm$0.43 -0.12$\pm$0.09 -0.30$\pm$0.15 -0.18$\pm$0.15 3.22$\pm$0.53
2 H 3.79$\pm$0.39 3.88$\pm$0.43 -0.15$\pm$0.11 -0.26$\pm$0.12 -0.26$\pm$0.19 3.16$\pm$0.58
S 3.64$\pm$0.39 3.64$\pm$0.29 -0.16$\pm$0.10 -0.27$\pm$0.16 -0.29$\pm$0.14 2.91$\pm$0.52
In our within-subject design study, the Shapiro-Wilk normality (SWN) test result shows that precision, recall and the final score data do not follow a normal distribution. Hence, we use the non-parametric Wilcoxon signed-rank (WSR) test. The results in Table VI show that there is no significant difference between captions generated by SoundSCaper and those offered by soundscape experts on the final score ($p=0.128$), and no significant difference between the two in terms of precision ($p$ = $0.34$) and recall ($p$ = $0.44$). This means that the quality of soundscape captions generated by SoundSCaper is comparable to that of soundscape expert-annotated captions. Fig. 8 details the precision, recall and final THumBS score in the evaluation of dataset D1. The horizontal line bisecting the box is the median, which coincides with the top line; the red dot represents the mean. The top and bottom borders of the box represent the 25th and 75th percentiles, respectively.
@@
Figure 8: P, R and THumBS score of soundscape captions given by a jury of 16 audio/soundscape experts in the human evaluation of dataset D1.
Figure 9: P, R and THumBS score of soundscape captions given by a jury of 16 audio/soundscape experts in the human evaluation of the external dataset D2.
2) RQ2: How do the captions from SoundSCaper and experts perform on the mixed external dataset D2?
@@
The SWN test results indicate that the final score data on D2 does not follow a normal distribution ($p<0.05$). Hence, we use a non-parametric WSR test. Table VI shows that expert-annotated captions scored slightly higher than SoundSCaper captions in human assessment; however, the WSR test result shows that there is no significant difference between the two on the final scores ($p=0.051$), as it close to the significant level. We evaluate the ratings on precision, recall and penalty items, including fluency, conciseness and irrelevance, respectively. The SWN test results in Fig. 9 show that the precision and recall ratings follow a normal distribution, while penalty items do not. Therefore, we use the paired t-test for precision and recall ratings; the result implies that there is no significant difference between the SoundSCaper and expert-annotated captions on precision rating ($p$ = $0.19$) while there is a significant difference in recall rating ($p$ = $0.028$), which is not surprising as SoundAQnet is untrained on those datasets and the AE labels are also limited. The WSR test result implies that there is no significant difference between the SoundSCaper and expert-annotated captions on fluency ($p$ = $0.33$), conciseness ($p$ = $0.97$), and irrelevance ($p$ = $0.21$). In summary, SoundSCaper has good generalization performance and adaptability, even though the recall rating of SoundSCaper captions is significantly lower than that of expert annotated captions, and a competitive final score is still achieved.
3) RQ3: On dataset D1, what are the cases with the biggest difference between captions from experts and SoundSCaper?
@@
The violin plot in Fig. 8 (c) reveals that the score distribution of SoundSCaper shows a slightly lower tail compared to that of the soundscape experts, indicating that SoundSCaper underperforms human experts on some audio clips. Here, we explore the largest gap in final scores between SoundSCaper and human experts by subtracting the final score of SoundSCaper captions from that of soundscape expert annotations. The maximum value in the difference sequence is 1.99, observed in the sample “28.flac”. The soundscape captions are:
Human soundscape expert: Immediately, there is a siren dominating the soundscape. As it fades away, it sounds like this is in a park, with people walking and chatting. Overall this is a calm soundscape, made somewhat annoying by the presence of the siren for part of the time.
SoundSCaper: In this park, you are surrounded by the soothing sounds of birds chirping, gentle human activities, and distant speech. The natural environment forms a tranquil backdrop. These sounds create a peaceful atmosphere, making you feel calm and content.
@@
In this case, where the expert caption outperforms SoundSCaper, it emphasizes the disruptive presence of a siren in the park scene, highlighting its significant impact on the soundscape’s calmness. Conversely, SoundSCaper paints a serene picture, not mentioning sirens and focusing only on peaceful elements like birds chirping and distant speech. The reason is that there are no sirens in the 15 classes of AE labels in the dataset used in this paper. As a result, the SoundAQnet of SoundSCaper fails to recognize AEs that it has not seen.
@@
Human soundscape expert: This is a busy urban square with a mix of sounds. There is a constant hubbub of people talking and light music. There are also light vehicles passing regularly. The character is pleasant, lively, and comfortable.
SoundSCaper: In this bustling public square, music fills the air, accompanied by the chatter of people and sounds of things clinking and rustling. Occasionally, the rumble of vehicles and traffic noise can be heard in the background. The atmosphere is lively and vibrant, making the scene eventful and far from monotonous.
@@
In this case, the SoundSCaper caption is better. The expert caption captures a mixture of pleasant, comfortable, and lively sounds in an urban square. SoundSCaper predominantly depicts the scene’s lively and vibrant aspects, such as music and people’s chatter, while downplaying vehicle noise. One possible explanation for this difference is that emotional feelings are subjective. People from various experiences and socio-cultural backgrounds may feel the same sound differently. Compared to descriptions with expert’s individual responses to AQs, the AQ values predicted by SoundAQnet, trained on the ARAUS dataset of 25248 samples assessed by 605 participants, may be less personalized and more acceptable to other participants.
4) RQ4: On dataset D2, what are the cases with the biggest difference between captions from experts and SoundSCaper?
Subtracting the final score of SoundSCaper captions from that of expert annotated captions, the maximum score difference is 1.89, the sample is “41.wav”, and the captions are:
Human soundscape expert: This urban street has a dense flow of vehicles passing from left to right and a stationary diesel engine dominating the right side of the soundscape. The character is eventful with moderate to high unpleasantness.
SoundSCaper: In this street, the predominant sound is that of vehicles passing by. It is neither calm nor pleasant, as it can be loud and overwhelming.
In this case, the expert caption outperforms SoundSCaper caption. The expert caption provides a detailed and contextual description, such as a dense flow of vehicles and a dominant stationary diesel engine, attributing a sense of unpleasantness to the scene. The SoundSCaper only emphasizes the overwhelming noise of passing vehicles without relevant contextual information. Compared with the delicate and thorough expert description, the SoundSCaper caption, which focuses on capturing the primary information, is broader.
Next, subtracting the final score of expert annotated captions from that of SoundSCaper, the maximum difference value is 2.12, the sample is “49.wav”, and the captions are:
Human soundscape expert: This highway has a slightly varying flow of vehicles passing from left to right. There is a very quiet siren off to the left. The sound is not particularly annoying since it is quiet, but it is unpleasant and fairly boring.
SoundSCaper: In this street, the predominant sound is the rumble of vehicles passing by. It lacks vibrant as the constant noise feels monotonous and can be overwhelming.
In this case, the SoundSCaper caption is better. The soundscape expert’s caption describes subtle sounds like a quiet siren and monotonous left-to-right traffic flow, while the SoundSCaper caption focuses on dominant sounds like the continuous rumbling of vehicles, describing the scene as monotonous and overwhelming. If there is a siren label in SoundAQnet’s training set, but for very quiet sirens, the empirical threshold we set at 0.3 may still remove them. Auditory perception varies between individuals, and human attention tends to favour dominant sounds influenced by primary AQs. Thus, some participants may not notice weak AEs like a quiet siren and be attracted by dominant sounds, and some evaluation experts commented that they could not hear the quiet siren, thus making SoundSCaper score higher than the expert in this case. Furthermore, similar to the case of “26.flac” in RQ3, the AQ values predicted by SoundAQnet, which is trained on the large-scale dataset, may be more consistent with most participants’ auditory and emotional experiences compared to expert captions with individual subjective responses to AQs.
V-D Discussion
In the case study of RQ3-4, human soundscape experts provided more specific and context-aware soundscape captions, e.g., they pointed out the direction of vehicle circulation and the presence of a stationary diesel engine, which increased the spatiality and realism of the soundscape. In contrast, SoundSCaper, which aims to describe the dominant sounds in soundscapes, generates general and broad captions, e.g., only the presence of vehicles is mentioned, and relevant detailed information is missing. SoundSCaper mentions dominant AEs but lacks information about the corresponding sound sources and their spatial and temporal distribution. Due to the limited AS and AE labels in the used dataset, the SoundAQnet cannot capture and focus on subtle but key sounds (such as short sirens) like soundscape experts in case of “28.flac”, further leading to the lack of detailed description capability of LLM for soundscapes. Soundscape experts’ captions are based on personal experiences and feelings with a certain subjective style, influenced by their personal experiences and their specialized training. However, AQs towards soundscapes vary among different individuals. The AQ predictor SoundAQnet is trained on a large-scale dataset with many participant reviews. Compared to the individual AQ responses of soundscape experts, the AQ values predicted by SoundAQnet have fewer personal characteristics and are more moderate and general, reducing the interpretation of personal subjective emotions. For example, soundscape experts may have a personal preference for AQs in their descriptions. In the case of “26.flac”, for the eventful soundscape, experts described it as comfortable and pleasant; in the case of “49.wav”, for the relatively quiet soundscape with car sounds, experts described it as fairly boring and unpleasant. In short, the performance of SoundSCaper in the human experts’ assessment is similar to that of the soundscape experts, and the difference between the two is insignificant. That is, SoundSCaper is competent for soundscape captioning in automated soundscape description.
VI Conclusion
Describing a soundscape, the perceived acoustic environment, is a cumbersome task. This paper formalizes the soundscape captioning task and proposes a deep-learning system, SoundSCaper, for this purpose. The SoundSCaper combines a dedicated acoustic model, SoundAQnet, with a large language model (LLM). Results show that the lightweight SoundAQnet can efficiently identify audio events (AEs) and acoustic scenes (ASs), and, most importantly, predict affective qualities (AQs). The last capability distinguishes our work from previous audio captioning efforts. Experiments prove that SoundAQnet relates human-perceived AQs with AEs and ASs in soundscapes, similarly as reported in the literature for humans.
A general LLM is then used to translate SoundAQnet’s predictions into natural language and add common knowledge on expectations for the detected specific AEs and contextual AS. To obtain high-quality captions, informative priming is used to make the LLM aware of the meaning of soundscape, while chain-of-thought prompting is used to ensure that all aspects of the soundscape are included in the caption. Based on comprehensive evaluation metrics, a jury of 16 audio/soundscape experts carefully evaluated the SoundSCaper-generated and the expert-annotated soundscape captions. The assessment results illustrate that the average scores (out of 5) of SoundSCaper-generated captions are lower than those of the experts by 0.21 and 0.25, respectively, but not statistically significant, on the evaluation set from the test set and on the external set consisting of 5 model-unknown datasets with varying lengths and acoustic properties. Overall, the captions generated by SoundSCaper perform close to those annotated by experts in human assessment. These findings suggest that the proposed automatic soundscape captioner system, SoundSCaper, can effectively describe audio clips from soundscapes automatically.
The next step will focus on the database limitations in this study, i.e., increasing the categories of audio events and scenes, and improving the combination and fusion way of the acoustic and language models in SoundSCaper to complete the end-to-end processing from audio information to natural language.

### v3

@@ V-E Correlation Study between AEs and AQs with SoundAQnet
Figure 7: The scatter plot, density distribution curves, and marginal distributions for the scenes involved are from SoundAQnet’s predictions on the test set.
V-F Analyzing ASs and AQs with SoundAQnet
@@
V-G SoundAQnet on the Cross-cultural Soundscape Dataset
The soundscape standard ISO/TS 12913-2 [35] gives researchers some freedom to incorporate cultural differences, but the basic idea behind the PAQ in [35] is still to try to remove cultural bias, starting with language. Soundscape attributes translation project (SATP) [56] aims at removing bias due to language and culture. The SATP dataset is a cross-cultural, multi-lingual, international, and cross-institutional dataset containing recordings with ISO-standard AQ labels. The SATP contains 19089 samples in 18 languages, consisting of 707 participants from 29 different national institutions evaluating 27 30-second binaural audio recordings according to ISO/TS 12913-2:2018 [35], in the institution’s own language in each region. There are an average of 32 participants per institution.
To evaluate SoundAQnet’s generalization ability across cultures, first, the Wilcoxon signed-rank test is used to compare the model’s predictions on SATP with the ratings of 707 participants. The results show no significant difference between SoundAQnet’s predictions and the average ratings of 707 participants ($p>0.05$); that is, SoundAQnet performs similarly to the participants. Then, to explore the model performance in specific cultures, the consistency analysis of two cases in Europe and Asia is performed using Intraclass Correlation Coefficient (ICC) [57] analysis. As shown in Table VII, SoundAQnet’s predictions show significant consistency with participants at University College London (UCL) on all AQs ($p<0.05$); and with participants at Nanyang Technological University (NTU) on 7 AQs ($p<0.05$) except Calm ($p>0.05$). This indicates that SoundAQnet shows significant consistency on multiple affective dimensions across two different cultures, including Asian and Western cultures.
TABLE VII: ICC analysis for two different cultures (* is $p<0.05$); ICC ranges from 0 (no consistency) to 1 (perfect consistency).
PAQ Plea Even Chao Vibr Unev Calm Anno Mono
UCL_ICC 0.89* 0.85* 0.89* 0.62* 0.82* 0.87* 0.82* 0.57*
NTU_ICC 0.78* 0.60* 0.90* 0.65* 0.61* 0.00 0.86* 0.53*
The results on SATP dataset show that SoundAQnet captures some degree of intra-cultural consistency. Prior studies show that LLMs generate different responses when queried in different languages, each embedding elements of its local culture [58]. Integrating SoundAQnet with LLMs may provide a good start for future studies aimed at fine-tuning the model to specific cultural contexts. In addition, we acknowledge that cultural differences in soundscapes are complex. Affective responses to soundscapes are inherently subjective and affected by personal, linguistic, and cultural factors, which is also one of the key challenges in soundscape studies.
VI Human evaluation (SoundSCaper experiment)
@@
VI-A Experimental Design for Quality Assessment
The study employs a within-subjects design to evaluate soundscape captions from SoundSCaper and human experts. The sample size calculation is performed using G*Power [59]. The results of the calculation indicated that a sample size of 30 audio samples with $\alpha=0.05$ and an assumed Effect Size of 0.5 for the Wilcoxon signed rank test achieved the pre-statistical power of 83.3%. Thus, the evaluation contains 60 audio clips from two distinct datasets. Dataset 1 (D1) contains 30 randomly selected samples with the same sound pressure levels (SPLs) from this paper’s test set. Dataset 2 (D2) has 30 samples randomly selected from 5 external, i.e., model-never-seen, audio scene datasets, which are DCASE 2018/2019 [13], ISD [30], LITIS-Rouen [60] and a road sound dataset [61] of freeways with extreme noise environments commonly seen in daily life. The training set used in this paper contains 3 types of acoustic scenes, so recordings related to the 3 AS labels are selected from the 5 external datasets. Finally, the total duration of the D2 candidate data pool is about 1177 hours. The audio clips in D2 vary from 10 to 30 seconds with various SPLs without any limitations. Hence, D2 is used mainly to test the generalization performance of the SoundSCaper system.
VI-A1 Soundscape expert annotations
@@
VI-A2 SoundSCaper captions
As described in Section IV-B, SoundSCaper automatically generates target descriptions.
Finally, 120 soundscape captions are evaluated, 60 of which are derived from the proposed SoundSCaper system, and the remaining 60 are annotated by the two experienced soundscape experts. These captions are randomized. These captions were evaluated by a jury of 16 audio/soundscape experts and another jury of 16 laypersons, totaling 32 participants from 8 countries, including the UK, Singapore, Belgium, Canada, and France. Human assessment instructions, assessment data, and statistics’ metadata are publicly available on the homepage. Based on a self-assessment of the study’s risks, ethical approval for this research was obtained from the Faculty of Engineering and Architecture of Ghent University.
VI-B Soundscape Caption Evaluation Metrics
Inspired by [62], we introduce the Transparent Human Benchmark for Soundscapes (THumBS) as a metric for the ASSC task. This indicator consists of precision, recall, and three other types of penalty items targeting specific defects.
VI-B1 Precision and recall $\in[1,5]$
@@ VI-B3 THumBS score
Due to limited space, we fully explain these metrics in the participant instructions presented on the homepage.
Figure 8: P, R and THumBS score of soundscape captions given by a jury of 16 audio/soundscape experts on the dataset D1.
Figure 9: P, R and THumBS score of soundscape captions given by a jury of 16 audio/soundscape experts on the external dataset D2.
VI-C Professionals’ Evaluations
1) Comparison of SoundSCaper and expert captions
TABLE VIII: Comparison of soundscape caption quality from soundscape expert (E) and SoundSCaper (S) on datasets D1 and D2.
@@ VI-C Professionals’ Evaluations
[…]
In the within-subject design study, the Shapiro-Wilk normality (SWN) test result shows that precision, recall and the final score data do not follow a normal distribution. Hence, we use the non-parametric Wilcoxon signed-rank (WSR) test. The results in Table VIII show that there is no significant difference between captions generated by SoundSCaper and those offered by soundscape experts on the final score ($p=0.128$), and no significant difference between the two in terms of precision ($p$ = $0.34$) and recall ($p$ = $0.44$). This means that the quality of soundscape captions generated by SoundSCaper is comparable to that of soundscape expert-annotated captions. Fig. 8 details the precision, recall and final THumBS score in the evaluation of dataset D1. The horizontal line bisecting the box is the median, which coincides with the top line; the red dot represents the mean. The top and bottom borders of the box represent the 25th and 75th percentiles, respectively.
@@
For the mixed external dataset D2, the SWN test results indicate that the final score data on D2 do not follow a normal distribution ($p<0.05$). Hence, we use a non-parametric WSR test. Table VIII shows that expert-annotated captions scored slightly higher than the captions generated by SoundSCaper; however, the WSR test shows that there is no significant difference between the two in the final scores ($p=0.051$). As $p$ is close to the significant level, we evaluate the ratings on the precision, recall, and penalty items, including fluency, conciseness, and irrelevance, separately. The distributions in Fig. 9 show that the precision and recall ratings follow a normal distribution, while the penalty items do not. Therefore, we use the paired t-test for precision and recall ratings; the result implies that there is no significant difference between the SoundSCaper-generated and expert-annotated captions on precision rating ($p$ = $0.19$) while there is a significant difference in recall rating ($p$ = $0.028$), which is not surprising as SoundAQnet is not trained on those datasets and the AE labels are also limited. The WSR test implies that there is no significant difference between SoundSCaper-generated and expert-annotated captions on fluency ($p$ = $0.33$), conciseness ($p$ = $0.97$), and irrelevance ($p$ = $0.21$). In summary, SoundSCaper has good generalization performance and adaptability, even though the recall rating of SoundSCaper-generated captions is significantly lower than that of the expert-annotated captions, and a competitive final score is still achieved.
2) Case study on the differences
@@
The violin plot in Fig. 8 (c) reveals that the score distribution of SoundSCaper shows a slightly lower tail compared to that of the soundscape experts, indicating that SoundSCaper underperforms on some audio clips. Here, we explore the largest gap in final scores between SoundSCaper and human experts by subtracting the final score of SoundSCaper captions from that of soundscape expert annotations. The maximum value in the difference sequence is 1.99, observed in the sample “28.flac”. The soundscape captions are:
@@
In this case, the expert outperforms SoundSCaper, and emphasizes the disruptive presence of a siren in the park scene, highlighting its significant impact on the soundscape’s calmness. Conversely, SoundSCaper paints a serene picture, not mentioning sirens and focusing only on peaceful elements such as birds chirping and distant speech. The reason is that there are no sirens in the 15 classes of AE labels in the dataset used for training the SoundAQnet, and hence it does not recognize this sound.
@@
In this case, the SoundSCaper caption is more appreciated by professionals. The expert’s caption captures a mixture of pleasant, comfortable, and lively sounds in an urban square. SoundSCaper predominantly depicts the scene’s lively and vibrant aspects, such as music and people’s chatter, while downplaying vehicle noise. One possible explanation for this difference is that emotional feelings are subjective. People from various experiences and socio-cultural backgrounds may feel the same sound differently. Compared to descriptions with experts’ individual responses to AQs, the AQ values predicted by SoundAQnet, trained on the ARAUS dataset of 25248 samples assessed by 605 participants, may be more consistent, less personalized, and more acceptable to other participants.
VI-D Layperson Evaluations
We employ the within-subject design for the layperson evaluation experiment. The SWN test shows that the final score data for D1 do not follow a normal distribution. Hence, we use the non-parametric WSR test. Table IX shows that there is no significant difference between captions generated by SoundSCaper and those provided by soundscape experts in terms of the final score ($p=0.136$) by the layperson jury. That is, in the opinion of laypersons, the quality of captions generated by SoundSCaper is comparable to that of soundscape expert-annotated captions.
For the external dataset D2, the SWN test results show that the final score data do not follow a normal distribution, so the non-parametric WSR test is used. Table IX shows that in the evaluation by a layperson jury, SoundSCaper captions scored 3.24, slightly higher than the expert-annotated captions. However, the WSR test results show no significant difference in the final score between the two ($p=0.90$).
TABLE IX: Layperson evaluation of soundscape caption from soundscape expert (E) and SoundSCaper (S) on datasets D1 and D2.
D precision recall fluency conciseness irrelevance Score
[…]
In addition, we compare the final scores given by 16 experts and 16 laypersons for both expert-annotated and SoundSCaper-generated captions. We use the paired t-test for expert ratings on both D1 and D2, as they follow the normal distribution. As for the final scores for SoundSCaper-generated captions, we employed the WSR test for both D1 and D2, as their ratings do not follow a normal distribution. The paired t-test results show no significant differences between the ratings of 16 experts and 16 laypersons for the expert-annotated captions on both D1 ($p=0.32$) and D2 ($p=0.54$). As for the final scores for SoundSCaper-generated captions, the WSR results show that there is no significant difference between the expert ratings and laypersons’ ratings ($p=0.58$) on D1; however, the laypersons’ ratings for SoundSCaper-generated captions are significantly higher than the expert ratings on D2 ($p<0.01$), laypersons rate the SoundSCaper-generated captions slightly higher than the expert annotations. This implies that laypersons are less sensitive to inconsistencies or subtle errors in soundscape captions compared to the experts, and appreciate the SoundSCaper-generated captions more.
VI-E Comparison with AAC systems
TABLE X: Average results of AAC systems and SoundScaper on various NLP metrics of the ASSC task on datasets D1 and D2, using soundscape expert captions as ground-truth sentences.
# System BLEU ROUGE-L METEOR CIDEr
Recall Precision F1 Score
1 P-LocalAFT 0.0249 0.1001 0.3428 0.1537 0.0684 0.0029
2 ConvNeXt-Trans 0.1115 0.1454 0.2542 0.1831 0.1168 0.0917
3 GAMA 0.0879 0.1280 0.2852 0.1694 0.1009 0.0849
4 Qwen-Audio 0.1032 0.1479 0.2729 0.1868 0.1082 0.1213
5 SoundSCaper 0.1901 0.2277 0.2131 0.2150 0.1610 0.2745
In this section, SoundSCaper is compared with AAC systems with and without LLMs. Table X shows the performance of typical sequence-to-sequence (S2S) systems without LLM trained on the AudioCaps and Clotho-v2 datasets: P-LocalAFT [23] and ConvNeXt-Trans [22], and audio-LLMs such as GAMA [24] and Qwen-Audio [25]. The audio encoder of ConvNeXt-Trans is ConvNeXt pre-trained on AudioSet [47], and the decoder consists of a Transformer decoder with a structure similar to GPT. Given the excellent performance of ConvNeXt-Trans on AAC tasks [22], we fine-tune ConvNeXt-Trans on the training set of this paper as the baseline system for soundscape audio-to-text with affective qualities. GAMA [24] uses the audio spectrum transformer (AST), which consists of 12 Transformer encoder layers, for audio encoding, and uses multiple models such as multi-layer aggregators and Q-Former for enhancement, and then uses GPT-4 for caption generation. The number of parameters and the number of multiply-add-accumulate operations (MACs) required for AST in GAMA are 73.61MB and 108.99 G, respectively, while the corresponding SoundAQnet in SoundSCaper is only 2.70MB and 27.68 G. In terms of both the number of parameters and the model complexity, the proposed acoustic model, SoundAQnet in SoundSCaper, is much smaller than AST in GAMA. In short, Table X shows that SoundSCaper performs better than the LLM-free S2S AAC systems and the audio-LLMs on several NLP metrics. Please note that the soundscape expert captions have no influence on SoundSCaper’s design, implementation, and training. In fact, they are collected after SoundSCaper is completed.
For source codes and data to fine-tune the audio-to-text baseline ConvNeXt-Trans for soundscapes, as well as scripts for NLP metrics, please see the homepage.
VI-F Further Discussions
In the case study in Section VI-C, soundscape experts provided more specific and context-aware soundscape captions; e.g., they pointed out that the siren dominates the soundscape first, and after it fades away, people are chatting, which increased the spatiality and realism of the soundscape. In contrast, SoundSCaper mentions dominant AEs but lacks information about the corresponding sound sources and their spatial and temporal distribution. Due to the limited AS and AE labels in the used dataset, SoundAQnet cannot capture subtle but key sounds (such as short sirens) on the case “28.flac” as soundscape experts, resulting in descriptions lacking details. In addition, individuals from different cultural backgrounds have different AQs for soundscapes. The descriptions of the two soundscape experts from the UK and Sweden are based on personal experience, perceptions, and specialized training, and have a certain subjective style. Compared with the individual AQ responses of soundscape experts, the AQ values predicted by SoundAQnet, which is trained based on 25248 samples assessed by 605 participants, are more agreeable to other assessors in the case “26.flac”. This is consistent with SoundAQnet’s cross-cultural performance on the SATP dataset, which contains 19089 samples in 18 languages and comprises 707 participants from 29 national institutions as discussed in Section V-G. In short, SoundSCaper’s acoustic model, SoundAQnet, can effectively predict soundscape AQ in human evaluation based on datasets D1 and external mixed D2 and shows certain generalization abilities on the SATP dataset across countries and cultures. However, SoundSCaper still faces challenges posed by the inherent bias and subjectivity of individuals’ affective response to sound and socio-cultural differences, which are key issues in soundscape research.
To assess the quality of ASSC, we investigated whether people would rate SoundScaper-generated and human-annotated captions differently after listening to soundscape recordings. In evaluations by 16 experts and 16 laypersons, captions produced by SoundSCaper are rated similarly to those produced as a consensus of two experts from different countries. The evaluation results show no significant difference in the rating behavior between the 16 experts and 16 laypersons. However, the 16 laypersons rated SoundSCaper higher on the D2 dataset, probably because laypersons are not as sensitive as experts to subtle differences and imprecision in soundscape captions.
In contrast to many AAC systems, SoundSCaper is not trained on datasets with captions. However, in Table X, where captions annotated by soundscape experts are used as ground truth, SoundSCaper outperforms AAC systems with and without LLM in NLP metrics. This shows that SoundSCaper outperforms general AAC systems in the ASSC task. However, since this study is the first attempt at ASSC, it faces some limitations: (1) the lack of datasets, preferably audiovisual datasets containing 360-degree videos and spatial audio, with high-quality soundscape captions produced by professionals; (2) the focus on common outdoor soundscapes, although the ARAUS dataset is based on USotW that contains recordings from 12 cities in 3 continents, it does not include extreme and rare soundscapes; (3) the subjectivity of individual perception of soundscapes and the differences in various regions and cultures need to be further explored. The proposed SoundSCaper’s generalization performance in practical applications still needs to be optimized with more diverse data to ensure its adaptability to a wider range of soundscapes and user groups.
VII Conclusion
Describing a soundscape, the acoustic environment as it is perceived and understood by people in context, is a cumbersome task. This paper formalizes the affective soundscape captioning (ASSC) task and designs the SoundSCaper system for this purpose. SoundSCaper used a novel lightweight acoustic model, SoundAQnet, to classify ASs and AEs and to predict AQs from soundscape, and also customized an LLM with prompt engineering techniques to turn such information into textual descriptions. Our work is significantly different from previous audio captioning studies, because it is the first attempt to jointly model the AS and AE in acoustic environments and the corresponding human-perceived AQ in soundscapes, and also the first attempt to connect the classification of AS and AE in acoustic environments with human affective descriptions of soundscapes. Comprehensive experiments and human evaluations have been performed to demonstrate the effectiveness of the proposed system for ASSC, as compared to relevant baselines. Overall, SoundSCaper offers competitive performance in human subjective evaluation and various objective captioning metrics, and the generated captions are comparable to those annotated by soundscape experts.
Given that this study is the first attempt at ASSC, it faces several limitations, such as the lack of available datasets with fully and carefully annotated soundscape recordings, and limited types of soundscapes and acoustic environments involved. In addition, AASC faces challenges posed by the inherent bias and subjectivity of individuals in their affective responses and socio-cultural differences, which are key issues in soundscape studies. Hence, the generalization of SoundSCaper needs to be further improved through diversified datasets and cross-scene and cultural validation. Introducing diverse large-scale data also helps the LLM in SoundSCaper fit AQs, explore scaling laws in AASC, and reduce the chance of LLM hallucinations, which remains a common issue. In addition, SoundAQnet performs comparably with human participants on the cross-country SATP dataset, implying that it has the potential to serve in unseen cross-cultural soundscape AQ assessments.